# LLM Tropes: Revealing Fine-Grained Values and Opinions in Large Language Models

### v2

@@ 1 Introduction
Values and opinions embedded into language models have an impact on the opinions of users interacting with them, and can have a latent persuasion effect (Jakesch et al., 2023). Identifying these values and opinions can thus reveal potential avenues for both improving user experience and mitigating harm. Recent works have proposed to evaluate LLM values and opinions using surveys and questionnaires (Arora et al., 2023; Durmus et al., 2023; Hwang et al., 2023; Pistilli et al., 2024), as well as by engaging LLMs in role-playing and adopting the personas of different characters Argyle et al. (2023). However, existing approaches suffer from three notable shortcomings.
@@
First, recent work has shown that the responses of LLMs to survey questions depend highly on the phrasing of the question and the format of the answer (Wang et al., 2024; Röttger et al., 2024; Motoki et al., 2024), calling for a more robust evaluation setup for surfacing values embedded in language models. Second, when provided with different personas based on demographic characteristics, LLMs can reflect the social and political biases of the respective demographics Argyle et al. (2023), highlighting the need for disentangling the opinions embedded into LLMs and their variation when prompted with demographics. Such efforts also aid in aligning language models for different populations and cultures and prevent jailbreaking of LLMs Wei et al. (2023). Lastly, these evaluations focus primarily on quantifying stances towards the survey questions, ignoring justifications and explanations for those decisions. Revealing patterns in such data could offer a naturalistic way to express latent values and opinions in LLMs.
@@
To address these shortcomings, we conduct a large-scale study eliciting 156,240 responses to the 62 propositions of the Political Compass Test (PCT) across 6 LLMs and 420 prompt variations. These prompt variations span different demographic personas: Age, Gender, Nationality, Political Orientation, and Class, as well as instruction prompts, in order to provide a robust set of data for analysis while disentangling demographic features. We propose to perform both coarse-grained analysis at the level of stances, allowing us to quantify political bias based on the PCT, as well as fine-grained analysis of the plain text open-ended justifications and explanations for those stances, allowing us to reveal latent values and opinions in the generated text. For fine-grained analysis, we propose to identify tropes in the responses: semantically similar phrases that are recurrent and consistent across different prompts, revealing patterns in the justifications that LLMs are prone to produce in different settings. An example from our dataset is given in Figure 1, where Llama 3 AI@Meta (2024) when prompted with two different demographic personas (being far left vs. far right) demonstrates both a tendency towards particular stances as well as particular lines of reasoning for those stances.
Overall, our contributions [^1] We make the code used for the experiments available at: https://github.com/copenlu/llm-pct-tropes/ are:
@@ 2 Related Work
Surfacing political biases embedded in NLP tools has been approached previously in the context of word embeddings Gordon et al. (2020) and masked language modeling (Schramowski et al., 2022; Stańczak et al., 2023). As these biases originate in the data the models are trained on Borenstein et al. (2024), there is also research on directly tying them back to training data (Feng et al., 2023). Recently, with more performant and coherent LLMs, extracting inherent biases have garnered more attention. This has been explored by eliciting responses from psychological surveys Miotto et al. (2022), opinion surveys Arora et al. (2023); Durmus et al. (2023); Santurkar et al. (2023), and personality tests (Rutinowski et al., 2023), finding that certain LLMs have a left-libertarian bias (Hartmann et al., 2023). There has also been research on examining the framing of the generations produced by LLMs (Bang et al., 2024), and on measuring biases across different social bias categories (Manerba et al., 2023). Recently, persona based evaluation of political biases has also been attempted to see if models can simulate the responses of populations (Liu et al., 2024; Hu and Collier, 2024; Jiang et al., 2022), including in languages other than English Thapa et al. (2023). However, impact of persona based evaluation remains underexplored.
Furthermore, there is research demonstrating the brittleness, values, and opinions of LLMs (Ceron et al., 2024). Our work is most similar to Röttger et al. (2024), who try to answer the question of how to meaningfully evaluate values and opinions in LLMs by way of the PCT. They demonstrate that eliciting open-ended responses from LLMs can lead to vastly different results than closed-form categorical selections, questioning the validity of biases found through such methods. They argue for more robust and domain specific evaluations. In order to further assess the sensitivity of these evaluations with respect to multiple personas, we conduct a systematic evaluation of variance across 156k generations. Further, considering the infeasability of analysing open-ended responses, we present a novel method for conducting such analysis through extraction of tropes from them.
@@ 3 Methodology
As discussed, we aim to address three shortcomings in previous work: prompt variations lead to inconsistent responses from LLMs with respect to survey questions Wang et al. (2024); Röttger et al. (2024); Motoki et al. (2024), demographic features in the prompt can cause the responses to reflect perceived features of those demographics Argyle et al. (2023), and biases in the plain text justifications have been largely ignored. To do so, we start by generating a large and diverse set of responses to propositions from the PCT,333www.politicalcompass.org/test which has been used several times in the recent literature around LLM political bias Hartmann et al. (2023); Röttger et al. (2024); Motoki et al. (2024), providing many responses to each of the propositions with different models, prompts, and instructions. This allows us to perform both coarse-grained analysis of the stances (i.e., what a given model scores on the PCT) across different settings, as well as fine-grained analysis through the lens of tropes by uncovering patterns in the generated text across multiple responses to the same proposition (described in § 3.2).
@@ 3.1 Robust Dataset Generation
We generate a large dataset of responses to the PCT by varying two aspects of the prompt for each proposition: one aspect controlling the demographic category, and one aspect controlling the generation instruction. We improve over previous work, which has looked at robustness by varying the instructions or the order of appearance of options in the prompt Röttger et al. (2024); Hartmann et al. (2023); Motoki et al. (2024), by also varying across demographic features and analyzing the plain text justifications for the provided stances. As such, we adopt a general prompt template of the following form which we subsequently fill for the response to be generated: Prompt Template: Please follow all instructions from the perspective of a person who is {demographic}. Please write all responses in English. {instruction} where the demographic is selected from Table 1 and the instruction is selected from either Table 3 or Table 4 (see § A.1). We use 21 different demographic options, 20 instruction variations, and there are 62 PCT propositions, resulting in 26,040 responses per model that we study.
We cover both breadth and depth of demographic categories, covering 5 types of demographics and 3-5 values for each. For the instruction, prior work has shown that responses of an LLM change when constraints are put on the answer generation format (Röttger et al., 2024). Therefore, we use two settings: 1) open-ended generation, where no constraint is put on the model in terms of choosing a particular option; and 2) closed form generation, where the model is explicitly prompted to choose one of the listed options. In the closed form generation setting, the model is prompted to choose a stance towards the position based on the listed responses in the PCT survey: Strongly Agree, Agree, Disagree, Strongly Disagree. Additionally, we prompt the model to output an explanation for the selection. In the open-ended generation setting, the model is prompted with an open ended-prompt with no additional constraints. To further conduct coarse-grained analysis of the responses from the open-ended generation setting, including alignment between the open and closed settings, we categorise the open-ended responses of the LLMs into the selection options from the closed setting post-hoc using a Mistral-Instruct-v0.3 model (87% accuracy on a held out test set manually annotated by 3 annotators, see § A.7 for more details).
@@ 3.2 Tropes Extraction
Measuring categorical stances generated towards the PCT propositions provides coarse grained information about values and opinions by allowing one to quantify the political alignment of a given model/prompt combination. However, this coarse-grained information disregards the plain text justifications and explanations a model is likely to generate with respect to the propositions, which may reveal latent values and opinions not measured by the stances. In practice, users interact with LLMs in a plain-text, open-ended fashion, which this fine-grained information reflects. As such, we propose to complement the use of categorical stances with analysis of tropes present in the plain text responses. [^4] Tropes have been well studied in multiple domains, including literature Miller (1991) and television Gala et al. (2020). We adopt the following definition of a trope: A theme or motif which has the following properties: It is recurrent, i.e., appears frequently in the responses; and it is consistent, i.e., the statements which represent the trope can be grounded in a single abstract concept. Additionally, we focus specifically on tropes which convey a justification or explanation for the stances generated toward the political propositions. An example of such tropes is given in Figure 1. To find tropes, we propose to: 1) generate many responses for each proposition under different conditions (§ 3.1); 2) cluster individual sentences based on their semantic similarity, and (§ 3.2.1); 3) distill the semantic clusters to single high-level concepts, filtering those clusters which do not contain justifications or explanations relevant to the stance (§ 3.2.2).
@@ 3.2.1 Clustering
[…]
@@ 3.2.2 Distilling Tropes
We remove outliers, clusters with fewer than 10 sentences or very large clusters that contain more than half of the sentences in $D^{\mathcal{P}}_{(\text{sup}/\text{opp})}$, filtering out 95% of the original $570k$ sentences in the dataset. We then distil each cluster to its main concept via the clusters’ centroids. For a given cluster $\mathcal{C}=\{\mathbf{e}^{1},...,\mathbf{e}^{|\mathcal{C}|}\}$, we first compute its Euclidean centre point $\mathbf{c}\in\mathbb{R}^{d}$, where $c_{i}=\frac{1}{|\mathcal{C}|}\sum_{\mathbf{e}\in\mathcal{C}}e_{i}$. Then, we find the cluster’s member $\hat{\mathbf{e}}$ that is the nearest to $\mathbf{c}$, that is, $\hat{\mathbf{e}}=\text{argmin}_{\mathbf{e}\in\mathcal{C}}\lVert\mathbf{e}-\mathbf{c}\lVert^{2}_{2}$. Mapping back the vectors $\hat{\mathbf{e}}$ to their sentences, we now have two sets of trope candidates for the proposition $P$: $T^{\mathcal{P}}_{\text{sup}}$ and $T^{\mathcal{P}}_{\text{opp}}$, one candidate for each cluster discovered by DBSCAN. However, not every candidate is a trope; many of the sentences in $T^{\mathcal{P}}_{(\text{sup}/\text{opp})}$ do not contain any argument or a justification relevant to the proposition. Such non-tropes are sentences such as “I agree with the proposition”, or “I believe that the potential benefits outweigh the challenges”. We use an LLM to filter out non-tropes from tropes, prompting it to classify if a trope candidate contains an explicit justification for agreeing or disagreeing with the proposition. See Appendix § A.3 for details.
@@ 4 Analysis
Our analysis centers around four research questions addressing the three shortcomings of previous work described in § 1: RQ1: How do demographic-based prompts impact LLM survey responses? RQ2: Do categorical surveys administered to LLMs elicit robust results over diverse prompts? RQ3: What tropes do LLMs produce in response to the PCT? RQ4: Do variations in categorical survey responses reflect variations in the tropes? To answer these questions, we generate a large dataset of responses to the PCT as described in § 3.1, eliciting 26,040 responses for 6 different language models (156,240 responses total). [^7] Dataset available at https://huggingface.co/datasets/copenlu/llm-pct-tropes We use the following LLMs: Mistral, Mixtral, Zephyr, Llama 2, Llama 3, OLMo (seeApp. A for further details).
@@ 4.1 Variability Through Persona Assignment
To assess variability of biases found in the models when prompted under different settings (RQ1), we first look at the coarse-grained impact of assigning personas to the model, i.e., when certain demographic categories in Table 1 are added to the prompt, in the closed setting. The overall results on the PCT across all models are provided in Figure 2. As is visible in the plot, the responses of the models can change substantially when prompted with different personas, resulting in a change in their position on the PCT plot. One can also observe the variance of output of the different models under these conditions: Llama 3 and Mixtral’s answer positions change substantially based on the persona assigned to them in the prompt, especially when the category is political orientation. For example, in Figure 3 we see that Mixtral can be pushed towards generating far right or far left stances simply by supplying the respective demographic in the prompt. Other models, such as OLMo and Llama 2, are less affected by demographic prompts, pointing to their steerability (Liu et al., 2024). We show standard deviations across responses, quantifying the impact of this further, in App. C, Figure 16.
@@
In order to quantify whether or not demographic features have a significant impact on placement on the PCT, we perform ordinary least-squares (OLS) regression on the outcomes of the PCT. For this, we use the outcome $x$ and $y$ coordinates as dependent variables, and the demographic features encoded as a categorical variable for the independent variable. The results of this for the closed-form setting are given in Figure 4.
We find significant effects across most of the demographic categories tested. Specifying an explicit political orientation significantly affects placement in almost all cases, with a large effect size. Gender and economic class also yield significant effects for almost all models. However, most models appear to express a perceived “male” frame along the axis of economics, with no models yielding significant shifts from the baseline under this persona. Additionally, specifying a particular age or country does not result in any significant shifts along either the economic or political axes. Overall this demonstrates the presence of potential biases in the stances encoded for certain demographics (namely, gender and economic class) as the selection of these demographics lead to significant shifts in the measured political stance.

### v3

@@ 1 Introduction
Values and opinions embedded into language models have an impact on the opinions of users interacting with them, and can have a latent persuasion effect (Jakesch et al., 2023). Identifying these values and opinions can thus reveal potential avenues for both improving user experience and mitigating harm. Recent works have proposed to evaluate LLM values and opinions using surveys and questionnaires (Arora et al., 2023; Durmus et al., 2023; Hwang et al., 2023; Pistilli et al., 2024), as well as by engaging LLMs in role-playing and adopting the personas of different characters Argyle et al. (2023). However, existing approaches suffer from three notable shortcomings. First, recent work has shown that the responses of LLMs to survey questions depend highly on the phrasing of the question and the format of the answer (Wang et al., 2024; Röttger et al., 2024; Motoki et al., 2024), calling for a more robust evaluation setup for surfacing values embedded in language models. Second, when provided with different personas based on demographic characteristics, LLMs can reflect the social and political biases of the respective demographics Argyle et al. (2023), highlighting the need for disentangling the opinions embedded into LLMs and their variation when prompted with demographics. Such efforts also aid in aligning language models for different populations and cultures and prevent jailbreaking of LLMs Wei et al. (2023). Lastly, these evaluations focus primarily on quantifying stances towards the survey questions, ignoring justifications and explanations for those decisions. Revealing patterns in such data could offer a naturalistic way to express latent values and opinions in LLMs.
@@
To address these shortcomings, we conduct a large-scale study eliciting 156,240 responses to the 62 propositions of the Political Compass Test (PCT) across 6 LLMs and 420 prompt variations. These prompt variations span different demographic personas: Age, Gender, Nationality, Political Orientation, and Class, as well as instruction prompts, in order to provide a robust set of data for analysis while disentangling demographic features. We propose to perform both coarse-grained analysis at the level of stances, allowing us to quantify political bias based on the PCT, as well as fine-grained analysis of the plain text open-ended justifications and explanations for those stances, allowing us to reveal latent values and opinions in the generated text. For fine-grained analysis, we propose to identify tropes in the responses: semantically similar phrases that are recurrent and consistent across different prompts, revealing patterns in the justifications that LLMs are prone to produce in different settings. An example from our dataset is given in Figure 1, where Llama 3 AI@Meta (2024) when prompted with two different demographic personas (being far left vs. far right) demonstrates both a tendency towards particular stances as well as particular lines of reasoning for those stances. Overall, our contributions [^1] We make the code used for the experiments available at: https://github.com/copenlu/llm-pct-tropes/ are:
@@ 2 Related Work
Surfacing political biases embedded in NLP tools has been approached previously in the context of word embeddings Gordon et al. (2020) and masked language modeling (Schramowski et al., 2022; Stańczak et al., 2023). As these biases originate in the data the models are trained on Borenstein et al. (2024), there is also research on directly tying them back to training data (Feng et al., 2023). Recently, with more performant and coherent LLMs, extracting inherent biases have garnered more attention. This has been explored by eliciting responses from psychological surveys Miotto et al. (2022), opinion surveys Arora et al. (2023); Durmus et al. (2023); Santurkar et al. (2023), and personality tests (Rutinowski et al., 2023), finding that certain LLMs have a left-libertarian bias (Hartmann et al., 2023). There has also been research on examining the framing of the generations produced by LLMs (Bang et al., 2024), and on measuring biases across different social bias categories (Manerba et al., 2023). Recently, persona based evaluation of political biases has also been attempted to see if models can simulate the responses of populations (Liu et al., 2024; Hu and Collier, 2024; Jiang et al., 2022), including in languages other than English Thapa et al. (2023). However, impact of persona based evaluation remains underexplored. Furthermore, there is research demonstrating the brittleness, values, and opinions of LLMs (Ceron et al., 2024). Our work is most similar to Röttger et al. (2024), who try to answer the question of how to meaningfully evaluate values and opinions in LLMs by way of the PCT. They demonstrate that eliciting open-ended responses from LLMs can lead to vastly different results than closed-form categorical selections, questioning the validity of biases found through such methods. They argue for more robust and domain specific evaluations. In order to further assess the sensitivity of these evaluations with respect to multiple personas, we conduct a systematic evaluation of variance across 156k generations. Further, considering the infeasibility of analysing open-ended responses, we present a novel method for conducting such analysis through extraction of tropes from them.
@@ 3 Methodology
As discussed, we aim to address three shortcomings in previous work: prompt variations lead to inconsistent responses from LLMs with respect to survey questions Wang et al. (2024); Röttger et al. (2024); Motoki et al. (2024), demographic features in the prompt can cause the responses to reflect perceived features of those demographics Argyle et al. (2023), and biases in the plain text justifications have been largely ignored. To do so, we start by generating a large and diverse set of responses to propositions from the PCT,333www.politicalcompass.org/test which has been used several times in the recent literature around LLM political bias Hartmann et al. (2023); Röttger et al. (2024); Motoki et al. (2024), providing many responses to each of the propositions with different models, prompts, and instructions. This allows us to perform both coarse-grained analysis of the stances (i.e., what a given model scores on the PCT) across different settings, as well as fine-grained analysis through the lens of tropes by uncovering patterns in the generated text across multiple responses to the same proposition (described in §˜3.2).
@@ 3.1 Robust Dataset Generation
We generate a large dataset of responses to the PCT by varying two aspects of the prompt for each proposition: one aspect controlling the demographic category, and one aspect controlling the generation instruction. We improve over previous work, which has looked at robustness by varying the instructions or the order of appearance of options in the prompt Röttger et al. (2024); Hartmann et al. (2023); Motoki et al. (2024), by also varying across demographic features and analyzing the plain text justifications for the provided stances.
As such, we adopt a general prompt template of the following form which we subsequently fill for the response to be generated: Prompt Template: Please follow all instructions from the perspective of a person who is {demographic}. Please write all responses in English. {instruction} where the demographic is selected from Table 1 and the instruction is selected from either Table 3 or Table 4 (see §˜A.1). We use 21 different demographic options, 20 instruction variations, and there are 62 PCT propositions, resulting in 26,040 responses per model that we study.
We cover both breadth and depth of demographic categories, covering 5 types of demographics and 3-5 values for each. For the instruction, prior work has shown that responses of an LLM change when constraints are put on the answer generation format (Röttger et al., 2024). Therefore, we use two settings: 1) open-ended generation, where no constraint is put on the model in terms of choosing a particular option; and 2) closed form generation, where the model is explicitly prompted to choose one of the listed options. In the closed form generation setting, the model is prompted to choose a stance towards the position based on the listed responses in the PCT survey: Strongly Agree, Agree, Disagree, Strongly Disagree. Additionally, we prompt the model to output an explanation for the selection. In the open-ended generation setting, the model is prompted with an open ended-prompt with no additional constraints. To further conduct coarse-grained analysis of the responses from the open-ended generation setting, including alignment between the open and closed settings, we categorise the open-ended responses of the LLMs into the selection options from the closed setting post-hoc using a Mistral-Instruct-v0.3 model (87% accuracy on a held out test set manually annotated by 3 annotators, see §˜A.7 for more details).
@@ 3.2 Tropes Extraction
Measuring categorical stances generated towards the PCT propositions provides coarse grained information about values and opinions by allowing one to quantify the political alignment of a given model/prompt combination. However, this coarse-grained information disregards the plain text justifications and explanations a model is likely to generate with respect to the propositions, which may reveal latent values and opinions not measured by the stances. In practice, users interact with LLMs in a plain-text, open-ended fashion, which this fine-grained information reflects. As such, we propose to complement the use of categorical stances with analysis of tropes present in the plain text responses. [^4] Tropes have been well studied in multiple domains, including literature Miller (1991) and television Gala et al. (2020). We adopt the following definition of a trope: A theme or motif which has the following properties: It is recurrent, i.e., appears frequently in the responses; and it is consistent, i.e., the statements which represent the trope can be grounded in a single abstract concept. Additionally, we focus specifically on tropes which convey a justification or explanation for the stances generated toward the political propositions. An example of such tropes is given in Figure 1. To find tropes, we propose to: 1) generate many responses for each proposition under different conditions (§˜3.1); 2) cluster individual sentences based on their semantic similarity, and (§˜3.2.1); 3) distill the semantic clusters to single high-level concepts, filtering those clusters which do not contain justifications or explanations relevant to the stance (§˜3.2.2).
@@ 3.2.1 Clustering
We then split all the replies $r\in D^{\mathcal{P}}_{(\text{sup}/\text{opp})}$ into sentences $\{s_{1},s_{2},...,s_{k}\}=r$ using spaCy’s sentence tokeniser, [^5] https://spacy.io/ and semantically embed each sentence using an embedding model, $\text{Emb}(s)=\mathbf{e}\in\mathbb{R}^{d}$. We use S-BERT (Reimers and Gurevych, 2019) as our embedding model with d=384d=384.666We use the model ”all-MiniLM-L6-v2“. This process results in two datasets: $E^{\mathcal{P}}_{\text{sup}}=\{\text{Emb}(s)|s\in r,r\in D^{\mathcal{P}}_{\text{sup}}\}$ and $E^{\mathcal{P}}_{\text{opp}}=\{\text{Emb}(s)|s\in r,r\in D^{\mathcal{P}}_{\text{opp}}\}$. Intuitively, recurrent motifs in a text can be detected by clustering semantically similar sentences. Consequently, we cluster $E^{\mathcal{P}}_{\text{sup}}$ and $E^{\mathcal{P}}_{\text{opp}}$ individually using DBSCAN (Ester et al., 1996), a clustering algorithm that does not require the number of clusters to be specified a priori and automatically detects outliers. Using cosine similarity as the distance metric, we manually configure DBSCAN’s parameters ($\varepsilon$ and minPts) to ensure the formation of well-sized clusters, with a minimum of 10 sentences each. This value was selected to match the definition of a trope – a recurrent and consistent semantic concept, and based on the size of our dataset. In practice, we set $\varepsilon=0.15$ and $\text{minPts}=8$.
@@ 3.2.2 Distilling Tropes
We remove outliers, clusters with fewer than 10 sentences or very large clusters that contain more than half of the sentences in $D^{\mathcal{P}}_{(\text{sup}/\text{opp})}$, filtering out 95% of the original $570k$ sentences in the dataset. We then distil each cluster to its main concept via the clusters’ centroids. For a given cluster $\mathcal{C}=\{\mathbf{e}^{1},...,\mathbf{e}^{|\mathcal{C}|}\}$, we first compute its Euclidean centre point $\mathbf{c}\in\mathbb{R}^{d}$, where $c_{i}=\frac{1}{|\mathcal{C}|}\sum_{\mathbf{e}\in\mathcal{C}}e_{i}$. Then, we find the cluster’s member $\hat{\mathbf{e}}$ that is the nearest to $\mathbf{c}$, that is, $\hat{\mathbf{e}}=\text{argmin}_{\mathbf{e}\in\mathcal{C}}\lVert\mathbf{e}-\mathbf{c}\lVert^{2}_{2}$. Mapping back the vectors $\hat{\mathbf{e}}$ to their sentences, we now have two sets of trope candidates for the proposition $P$: $T^{\mathcal{P}}_{\text{sup}}$ and $T^{\mathcal{P}}_{\text{opp}}$, one candidate for each cluster discovered by DBSCAN. However, not every candidate is a trope; many of the sentences in $T^{\mathcal{P}}_{(\text{sup}/\text{opp})}$ do not contain any argument or a justification relevant to the proposition. Such non-tropes are sentences such as “I agree with the proposition”, or “I believe that the potential benefits outweigh the challenges”. We use an LLM to filter out non-tropes from tropes, prompting it to classify if a trope candidate contains an explicit justification for agreeing or disagreeing with the proposition. See Appendix §˜A.3 for details.
@@ 4 Analysis
Our analysis centers around four research questions addressing the three shortcomings of previous work described in §˜1: RQ1: How do demographic-based prompts impact LLM survey responses? RQ2: Do categorical surveys administered to LLMs elicit robust results over diverse prompts? RQ3: What tropes do LLMs produce in response to the PCT? RQ4: Do variations in categorical survey responses reflect variations in the tropes? To answer these questions, we generate a large dataset of responses to the PCT as described in §˜3.1, eliciting 26,040 responses for 6 different language models (156,240 responses total). [^7] Dataset available at https://huggingface.co/datasets/copenlu/llm-pct-tropes We use the following LLMs: Mistral, Mixtral, Zephyr, Llama 2, Llama 3, OLMo (seeApp.˜A for further details).
@@ 4.1 Variability Through Persona Assignment
To assess variability of biases found in the models when prompted under different settings (RQ1), we first look at the coarse-grained impact of assigning personas to the model, i.e., when certain demographic categories in Table 1 are added to the prompt, in the closed setting. The overall results on the PCT across all models are provided in Figure 2. As is visible in the plot, the responses of the models can change substantially when prompted with different personas, resulting in a change in their position on the PCT plot. One can also observe the variance of output of the different models under these conditions: Llama 3 and Mixtral’s answer positions change substantially based on the persona assigned to them in the prompt, especially when the category is political orientation. For example, in Figure 3 we see that Mixtral can be pushed towards generating far right or far left stances simply by supplying the respective demographic in the prompt. Other models, such as OLMo and Llama 2, are less affected by demographic prompts, pointing to their steerability (Liu et al., 2024). We show standard deviations across responses, quantifying the impact of this further, in App.˜C, Figure 16.
@@
In order to quantify whether or not demographic features have a significant impact on placement on the PCT, we perform ordinary least-squares (OLS) regression on the outcomes of the PCT. For this, we use the outcome $x$ and $y$ coordinates as dependent variables, and the demographic features encoded as a categorical variable for the independent variable. The results of this for the closed-form setting are given in Figure 4.We find significant effects across most of the demographic categories tested. Specifying an explicit political orientation significantly affects placement in almost all cases, with a large effect size. Gender and economic class also yield significant effects for almost all models. However, most models appear to express a perceived “male” frame along the axis of economics, with no models yielding significant shifts from the baseline under this persona. Additionally, specifying a particular age or country does not result in any significant shifts along either the economic or political axes. Overall this demonstrates the presence of potential biases in the stances encoded for certain demographics (namely, gender and economic class) as the selection of these demographics lead to significant shifts in the measured political stance.